OpenStudy (anonymous):

the ages of elsa, edna and ellie are consecutive intergers. the sum of their ages is 126. what are their ages?

hartnn (hartnn):

let the age of elsa be x since the ages are consecutive, ages of other 2 are x+1 and x+2 sum = 108 so x+x+1+x+2 = 108 can u solve further ?

OpenStudy (anonymous):

sory not 108 its 126

hartnn (hartnn):

oops, so x+x+1+x+2 = 126 can u find x ?

hartnn (hartnn):

@qnqwiz did u get what i have done ?

OpenStudy (anonymous):

im still working on it

OpenStudy (anonymous):

im sooooo confused..if the sum of their ages is 114 instead of 126 for their age..how do i solve for x if i dont know what im even looking for? i got 39 40 and 41....i dont know what im doing wrong!!:(

OpenStudy (anonymous):

the new some is 114 becuase i got the 126 problem wrong and they gave me a different problem

hartnn (hartnn):

now 114 ? x+x+1+x+2 = 114 combine like terms x+x+x=3x 1+2=3 so 3x+3=114 3x=111 x=37 got this? so ages are 37,38,39

OpenStudy (anonymous):

if the previous sum was 114, why do i get 111 though?

hartnn (hartnn):

because its \(\huge3x+3=114\) so when u subtract 3 from both sides, u get \(\huge3x+3-3=114-3=111\)

OpenStudy (anonymous):

ooh ok now i get it!!! thank you

hartnn (hartnn):

welcome :)